# Feature Distillation is the Better Choice for Model-Heterogeneous Federated Learning

### v2

@@ 1 Introduction
Federated Learning (FL) has become a core method for collaboratively training neural networks across distributed clients while ensuring data privacy is protected mcmahan2017communication ; wang2023fedcda ; DBLP:conf/cvpr/WangCW022 . Recently, this framework has attracted considerable attention and is being applied in various domains, such as autonomous driving systems li2021privacy ; nguyen2022federated and intelligent healthcare solutions dong_tpami ; pfitzner2021federated .
Typically, FL has been actively studied with a homogeneous model setting. With the development of IoT products, different devices often possess varying computing resources, namely different model training capabilities pfeiffer2023federated . In such a scenario, it is essential to explore model-heterogeneous federated learning (Hetero-FL) to maximize the utilization of distributed computing resources. The key challenge here lies in how to aggregate the shared knowledge from different models.
@@ 2 Related Work
Federated Learning is a framework that trains a unified global model by aggregating individual models from multiple clients, each trained on their locally stored datasets mcmahan2017communication ; DBLP:conf/cvpr/WangCW022 ; wang2023fedcda . One of the notable FL architectures is FedAvg mcmahan2017communication , which strengthens the global model by aggregating parameters from locally trained models on private data. li2020federated involves incorporating a proximal term to mitigate the impact of differing data distributions across devices. While these methods are devoted to the homogeneous model across clients, diao2020heterofl proposes to enable the training of heterogeneous local models with varying computation complexities on different clients. Similar to diao2020heterofl , wang2023flexifed integrates flexibility in both width and depth, utilizing skip connections to bypass certain layers and structured pruning to manage width. kang2023nefl is a strategy that adjusts to varying depths by aggregating common layers from clients’ networks, such as the VGG network, to create global models from different layer groups. This paper focuses on federated learning with heterogeneous models across clients with improved knowledge distillation techniques.

### v3

@@ 1 Introduction
Federated Learning (FL) has become a core method for collaboratively training neural networks across distributed clients while ensuring data privacy is protected mcmahan2017communication ; wang2023fedcda ; li2024unleashing . Recently, this framework has attracted considerable attention and is being applied in various domains, such as autonomous driving systems li2021privacy ; nguyen2022federated , recommender systems li2025www ; li2025systematic , and intelligent healthcare solutions dong_tpami ; pfitzner2021federated .
Typically, FL has been actively studied with a homogeneous model setting. With the development of IoT products, different devices often possess varying computing resources, namely different model training capabilities pfeiffer2023federated ; li2024pfeddil . In such a scenario, it is essential to explore model-heterogeneous federated learning (Hetero-FL) to maximize the utilization of distributed computing resources. The key challenge here lies in how to aggregate the shared knowledge from different models.
@@ 2 Related Work
Federated Learning is a framework that trains a unified global model by aggregating individual models from multiple clients, each trained on their locally stored datasets Li_2024_CVPR ; DBLP:conf/cvpr/WangCW022 ; li2025re ; lifedssi . One of the notable FL architectures is FedAvg mcmahan2017communication , which strengthens the global model by aggregating parameters from locally trained models on private data. li2020federated involves incorporating a proximal term to mitigate the impact of differing data distributions across devices. While these methods are devoted to the homogeneous model across clients, diao2020heterofl proposes to enable the training of heterogeneous local models with varying computation complexities on different clients. Similar to diao2020heterofl , wang2023flexifed integrates flexibility in both width and depth, utilizing skip connections to bypass certain layers and structured pruning to manage width. kang2023nefl is a strategy that adjusts to varying depths by aggregating common layers from clients’ networks, such as the VGG network, to create global models from different layer groups. This paper focuses on federated learning with heterogeneous models across clients with improved knowledge distillation techniques.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This work is supported by the National Key Research and Development Program of China under grant 2024YFC3307900; the National Natural Science Foundation of China under grants 62376103, 62302184, 62436003 and 62206102; Major Science and Technology Project of Hubei Province under grant 2024BAA008; Hubei Science and Technology Talent Service Project under grant 2024DJC078; Ant Group through CCF-Ant Research Fund; and Fundamental Research Funds for the Central Universities under grant YCJJ20252319. The computation is completed in the HPC Platform of Huazhong University of Science and Technology.